OpenStudy (ramennoodle01):

Find the zeros of this function f (x)=x^2-25

OpenStudy (anonymous):

You are just being asked to treat the function as a quadratic equation and solve for \[x\]

OpenStudy (anonymous):

that is it

OpenStudy (anonymous):

Actually thats just factoring, you have to find the values of \[x\] that 'zero' the function.

OpenStudy (jonnyvonny):

True, so, after you factor, get x alone for each x in the "()" i.e. (x-5)(x+5) x-5=0 x=5 I think you can get the other one

OpenStudy (anonymous):

\[ x^2-25\qquad\text{is of the form}\\ a^2-b^2=(a+b)(a-b)\\\text{we can write}\\ x^2-25^2=(x)^2-(5)^2\\=(x+5)(x-5) \\\text{to find roots, set it equal to } 0\\ (x+5)(x-5)=0\\\text{i.e., either}\\ x+5=0\qquad{\rm or}\qquad x-5=0\\ \implies \boxed{x=-5}\qquad{\rm or}\qquad \boxed{x=5} \]

OpenStudy (anonymous):

kapeesh?

OpenStudy (jonnyvonny):

At least let the man think, @electrokid .